Find the equation of the line that is perpendicular to the line -x+3y=-1 and goes through the point (4,-2)
Answer by lwsshak3(11628)

Find the equation of the line that is perpendicular to the line -x+3y=-1 and goes through the point (4,-2)
***
Standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
-x+3y=-1
3y=x-1
y=x/3-1/3
slope=1/3
..
slope of perpendicular line=-3 (negative reciprocal)
equation: y=-3x+b
solve for b using coordinates of given point (4,-2) on the line
-2=-3*4+b
b=10
equation: y=-3x+10